One angle measures 54 more degrees than 3 times another angle. The angles are supplementary. What are their measures?
Answer by fractalier(6550)

Call the angles x and y.
We have x + y = 180 and
x = 54 + 3y
If we substitute we get
54+3y + y = 180
4y + 54 = 180
4y = 126
y = 31.5 degrees
so that
x = 148.5 degrees